OpenStudy (anonymous):

log3 8........ how would i solve this by using a change of base? i got, log3 8=x 3^x=8

OpenStudy (arbershabani97):

you should continue adding the natural or normal logs like ln3^x=ln8 xln3=ln8 x=ln8/ln3, ln cannot be simplified x=1.892789261

OpenStudy (anonymous):

your the best :)

OpenStudy (arbershabani97):

thanks :P